# Irreducible finite-index representations of automorphism groups of trees

### v1

@@ Abstract
We consider a locally finite tree $T$ where every vertex has degree 3$\geq$ 3≥ 3, and a closed noncompact subgroup $G$ of the automorphism group $\mathrm{Aut}(T)$, with Tits’ independence property (P) and with transitive action on the boundary $\partial T$. We show that $G$ does not have continuous irreducible representations to a locally convex topological vector space over $\mathbb{R}$, $\mathbb{C}$ or $\mathbb{H}$ that preserve a continuous nondegenerate sesquilinear form of index $p>1$. As the representations of index 111 are already classified, this gives a complete classification of continuous irreducible representations of $G$ that preserve a continuous nondegenerate sesquilinear form of finite index.
@@ 1 Introduction
In this paper, we study automorphism groups of locally finite trees and their continuous irreducible linear representations. We focus on a class of groups that satisfy Tits’ independence property (P) ([19], 4.2) and act transitively on the boundary at infinity of the tree. We assume that every vertex of the tree has degree 3$\geq$ 3≥ 3, in which case the hypotheses imply that the tree is homogeneous or semi-homogeneous. The groups we consider include universal groups associated with 222-transitive permutation groups, which were introduced by Burger and Mozes in ([4], 3.2) and then studied by Amann in his Ph.D. thesis ([1], Chapter 3). More recently, Reid and Smith [18] provided a characterization of tree automorphism groups with Tits’ independence property (P) in terms of combinatorial structures called local action diagrams.
In his Ph.D. thesis ([1], Chapter 2), Amann has classified continuous irreducible unitary representations of these groups, which fall into three families: super-cuspidal, special and spherical, according to the conjugacy class of a maximal compact subgroup that fixes a nontrivial vector. Our goal here is to classify continuous irreducible representations that are not unitary but preserve a continuous nondegenerate sesquilinear form of finite index instead. From now on, all sesquilinear from will be implicitly assumed to be continuous.
In general, we will work on a locally compact topological vector space over the field of real, complex, or quaternionic numbers. The existence of a sesquilinear form implies that the space has a pre-Hilbert structure, but we do not assume completeness. Irreducible representations are meant in the algebraic sense: the representation does not preserve any proper vector subspace except the trivial one. Our representations can also be regarded as continuous homomorphisms of $G<\mathrm{Aut}(T)$, where $T$ is a homogeneous or semi-homogeneous tree and $G$ satisfies the above assumptions, to $\mathrm{O}(p,\infty)$, $\mathrm{U}(p,\infty)$, or $\mathrm{Sp}(p,\infty)$, with $p\geq 1$.
@@
The study of Hilbert spaces with indefinite bilinear forms dates back to Pontrjagin [17], Iohvidov-Krein [10], and Naimark [13, 15, 12, 14]. More recently, Gromov ([9], Section 6) has suggested the study of the infinite-dimensional symmetric spaces of finite rank that can be defined subsequently. The case of the infinite-dimensional real hyperbolic space has been studied by Burger, Iozzi and Monod in [3], and later by Monod and Py in [11], while the general case has been studied by Duchesne in recent times [5, 8, 7, 6]. It turns out ([7], Corollary 1.10) that the only irreducible infinite-dimensional symmetric spaces of non-positive curvature and finite rank are, up to homothety, $\mathrm{O}(p,\infty)/\mathrm{O}(p)\times\mathrm{O}(\infty)$, $\mathrm{U}(p,\infty)/\mathrm{U}(p)\times\mathrm{U}(\infty)$, and $\mathrm{Sp}(p,\infty)/\mathrm{Sp}(p)\times\mathrm{Sp}(\infty)$, and that their isometry groups are, up to a subgroup of index two in the complex case, exactly the projective versions of $\mathrm{O}(p,\infty)$, $\mathrm{U}(p,\infty)$, and $\mathrm{Sp}(p,\infty)$ ([6], Theorem 3.3). Therefore, studying representations of $G$ that preserve a nondegenerate sesquilinear form of finite index is the same as studying actions of $G$ by isometries on infinite-dimensional symmetric spaces of nonpositive curvature and finite rank.
@@
In another paper [20], we consider the case where $G$ is a higher-rank algebraic group over a non-archimedean local field $\mathbb{F}$, and we show that, at least in the case $G=\mathrm{SL}_{n}(\mathbb{F})$ ($n\geq 3$), there are no continuous fixed point-free actions of $G$ on any infinite-dimensional symmetric space of nonpositive curvature and finite rank, which implies that there are no continuous irreducible representations of $G$ that preserve a nondegenerate sesquilinear form of finite index.
Turning back to the case where $G$ is a subgroup of the automorphism group of a tree, continuous irreducible representations to $\mathrm{O}(1,\infty)$ correspond to continuous actions by isometries on the infinite-dimensional real hyperbolic space, and these have been classified by Burger, Iozzi and Monod ([3], Theorem C), at least in the case where the tree $T$ is homogeneous and the action of $G$ is triply transitive on the boundary $\partial T$. The classification consists of a family of representations that depend on a single real parameter.
Our main result in this paper is that the classification by Burger, Iozzi and Monod basically exhausts all possible continuous irreducible representations of finite index. In fact we will show that, if $p>1$, there exist no continuous irreducible representations of $G$ to $\mathrm{O}(p,\infty)$, $\mathrm{U}(p,\infty)$, or $\mathrm{Sp}(p,\infty)$, and that in the case $p=1$ all complex and quaternionic representations are obtained by taking the tensor product with $\mathbb{C}$ or $\mathbb{H}$ of a real representation. The proof is done by generalizing the classification carried out by Amann ([1], Chapter 2) to the case where the representations are not necessarily unitary. Our classification is a generalization of the one carried out by Ol’shanskii [16] for $G=\mathrm{Aut}(T)$, except that we do not dive into the details of unitary representations but focus on those that can preserve indefinite sesquilinear forms.
@@
The paper is organized as follows. In Section 2 we give a few preliminaries on trees and on the independence property, which allow us to precisely state our hypotheses and the thesis of our theorem. In Section 3 we present the classification of the representations as spherical, special, and super-cuspidal, and we show that for our purposes it suffices to consider spherical representations. In Section 4 we consider the case where $G$ acts transitively on the set of vertices of the tree and give an explicit description of spherical representations, getting a classification that depends on only one parameter. In Section 5 we study nondegenerate sesquilinear forms of finite index that can be preserved by these representations, completing the proof of our theorem in the case where the action of $G$ is transitive on the vertices. Finally, in Section 6 we generalize the results of Sections 4 and 5 to the case where the action of $G$ is not transitive on the vertices.
@@ 2 Preliminaries
[…]
@@
We consider a locally finite tree $T$ where every vertex has degree 3$\geq$ 3≥ 3, and its automorphism group $\mathrm{Aut}(T)$. This group, equipped with the topology of pointwise convergence, is totally disconnected and locally compact. It has a natural action on the tree and on its boundary at infinity.
@@
We are ready to introduce the independence property for a subgroup $G<\mathrm{Aut}(T)$. We use the definition given by Amann ([1], Definition 9):
@@
As observed by Amann (as a consequence of Lemma 10 in [1]), if $G$ is a closed subgroup of $\mathrm{Aut}(T)$, this property is equivalent to Tits’ Property (P), which was defined in ([19], 4.2).
We are interested in studying representations of closed noncompact subgroups of $\mathrm{Aut(T)}$ that act transitively on the boundary $\partial T$ and have the independence property. From now on, let $G$ be such a subgroup. We recall a few properties of $G$ that follow from the assumptions; more details can be found in [1] (Proposition 4).
@@
    Either the action of $G$ on $V$ is transitive or it has two orbits. In the first case, the tree is homogeneous; in the second case, the tree is homogeneous or semi-homogeneous, one orbit is the set of vertices at even distance from a fixed vertex $o$, and the other orbit is the set of vertices at odd distance from $o$.
We are ready to state our main theorem. We use the notation $(\pi,H)$ for representations of $G$, where $H$ is a locally convex topological vector space over $\mathbb{K}\in\{\mathbb{R},\mathbb{C},\mathbb{H}\}$, and $\pi:G\rightarrow\mathrm{Aut}(H)$ is a group homomorphism. To avoid ambiguity, in the case $\mathbb{K}=\mathbb{H}$ we say that we consider a right vector space (i. e. where the scalar multiplication is on the right), so that actions are on the left.
Theorem 2.3.
Let $T$ be a locally finite tree where every vertex has degree 3$\geq$ 3≥ 3. Let $G$ be a closed noncompact subgroup of $\mathrm{Aut}(T)$, with the independence property and transitive action on $\partial T$. Let $(\pi,H)$ be a continuous irreducible $\mathbb{K}$-representation of $G$ to a locally convex topological vector space. Suppose that there exists a continuous nondegenerate finite-index sesquilinear form $B$ on $H$ that is preserved by $\pi$. Then the index of $B$ is 0 or 111. Moreover, if the index is 111, then the representation $(\pi,H)$ is isomorphic to the tensor product with $\mathbb{K}$ of a real representation.
We start with a simple but important observation.
@@
Let $(\pi,H)$ be a continuous irreducible representation of $G$. Then $(\pi,H)$ is smooth, i. e. every $v\in H$ is fixed by some open compact subgroup of $G$.
@@
Since the representation is irreducible and the subspace of smooth vectors, i. e. vectors that are fixed by some open compact subgroup of $G$, is $G$-invariant (if $g\in G$ and $K<G$ satisfies $Kv=v$, then $gv$ is fixed by the subgroup $gKg^{-1}$), it is enough to show that there exists one nonzero smooth vector.
[…]
@@ 3 Restricting to spherical representations
In this section, we begin the classification of the representations. First, we define:
@@
Given a continuous representation $(\pi,H)$ of $G$, we know from Proposition 2.4 that there is some open compact subgroup $K<G$ such that $H^{K}\neq\{0\}$, where the notation indicates the subspace of $K$-fixed vectors in $H$. We can consider the set of complete subtrees $S$ such that $H^{G(S)}\neq\{0\}$. What we have said, together with Proposition 2.2(a), implies that this set is not empty. Thus, we can consider the set $M$ of minimal complete subtrees $S$ in the set and distinguish three cases:
@@
Consider the contragredient representation $(\widetilde{\pi},\widetilde{H})$ of $(\pi,H)$. Since $H^{G(S)}\neq\{0\}$, we have $\widetilde{H}^{G(S)}\neq\{0\}$: the space $H^{G(S)}$ has a $G(S)$-invariant complement in $H$ and hence is isomorphic to $\widetilde{H}^{G(S)}$ ([2], Proposition 2.7).
@@
is a well-defined map from $H$ to $C(S)$: the function $g\mapsto\left\langle\xi,\pi(g^{-1})v\right\rangle$ is right-invariant under $G(S)$ because $\xi$ is invariant under $G(S)$, and left-invariant under some open compact subgroup of $G$ because every $v\in H$ is smooth. It is immediate to see that the map is $G$-equivariant and nonzero. Therefore, since the representation $(\pi,H)$ is irreducible, the map is injective and $(\pi,H)$ is a subrepresentation of $(\Pi_{S},C(S))$.
Amann proves ([1], Lemma 20) that if the representation $(\pi,H)$ is super-cuspidal, the space $H\subset C(S)$ that arises by applying Proposition 3.2 to $S\in M$ consists of functions that are finitely supported when projected to $G/S$. This is shown for complex representations, but the proof is the same for real and quaternionic ones. Amann then proceeds to classify unitary super-cuspidal representations, but the proof that the functions are finitely supported on $G/S$ works without assuming that the representation is unitary. This follows as a consequence, as the representation will preserve the scalar product on $H$ defined as the restriction to $H$ of the standard scalar product on the space of finitely supported functions on $G/S$.
Similarly, Amann proves ([1], Lemma 34) that if the representation $(\pi,H)$ is special, all functions in $H\subset C(S)$ (with $S$ consisting of two vertices linked by an edge) belong to $\ell^{2}(G/S)$. Again, the proof does not need that the representation is unitary, and this fact follows as a consequence, as the restriction to $H$ of the standard scalar product on $\ell^{2}(G/S)$ will be preserved.
In the rest of the paper, we will focus our attention on the remaining case, the case of spherical representations. Since in the other two cases all irreducible representations are unitary and since they satisfy $\dim(H^{G(S)})<\infty$ ([1], Lemma 20, Lemma 34), they cannot preserve a nondegenerate sesquilinear form of index 1$\geq$ 1≥ 1. This is explained in the following proposition.

### v2

# Irreducible representations of tree automorphism groups into Pontryagin spaces

@@ Abstract
Let $G=\mathrm{Aut}(T)$ be the automorphism group of a regular tree $T$. We study continuous irreducible representations of $G$ that preserve a continuous strongly nondegenerate sesquilinear form of finite index on a Hilbert space. These are already classified for index 0 (unitary case) and for index $1$. We show that there are no more representations for index >1>1, which completes the classification.
@@ 1 Introduction
Let $G=\mathrm{Aut}(T)$ be the automorphism group of a regular tree $T$, i. e. a tree where all vertices have the same degree $d\geq 3$. Irreducible algebraic representations of $G$ have been first studied and classified by Ol’shanskii [undefz]. Later, a complete and detailed classification of unitary representations of $G$ was provided by Figa-Talamanca and Nebbia [undefn].
Beyond the unitary world, a focus of much recent interest is representations preseving a sesquilinear form of finite index, or representations into Pontryagin spaces, first introduced in [undefab]. The unitary case corresponds to index 0. All index $1$ representations of $G$ have been classified by Burger, Iozzi, and Monod [undefe].
The main result of this paper is that G does not admit continuous irreducible representations of index >1>1. This is in strong contrast to the case of $\mathrm{SL}_{2}(\mathbb{R})$ and other rank one Lie groups, which admit representations of arbitrarily high index [undefh, undefu].
Theorem 1.1.
Let $G=\mathrm{Aut}(T)$, where $T$ is a locally finite regular tree. Let $\pi$ be a continuous irreducible representation of $G$ into a Hilbert space $H$. If $\pi$ preserves a continuous strongly nondegenerate sesquilinear form of finite index $p\geq 0$, then $p=0$ or $p=1$.
Strongly nondegenerate sesquilinear forms were defined in [undefe] as being associated with positive definite forms that generate a complete Hilbert norm. Traditionally, Hilbert spaces equipped with a strongly nondegenerate sesquilinear form are called Krein spaces, and Pontryagin spaces if the sesquilinear form has finite index.
Together with [undefz, undefn, undefe], Theorem 1.1 completes the classification of all continuous irreducible representations of $G$ to all Pontryagin spaces.
@@
The study of Hilbert spaces with indefinite sesquilinear forms dates back to Pontryagin [undefab], Iohvidov-Krein [undefp], and Naimark [undefw, undefy, undefv, undefx]; further treatment can be found in [undefa, undefc, undefq]. More recently, Gromov ([undefo], Section 6) has suggested the study of the infinite-dimensional symmetric spaces of finite rank that can be defined subsequently. The case of the infinite-dimensional real hyperbolic space has been studied by Burger, Iozzi, and Monod in [undefe], and later by Monod and Py in [undefu], while the general case has been studied by Duchesne in recent times [undefj, undefm, undefl, undefk]. It turns out ([undefl], Corollary 1.10) that the only irreducible infinite-dimensional symmetric spaces of non-positive curvature and finite rank are, up to homothety, $\mathrm{O}(p,\infty)/\mathrm{O}(p)\times\mathrm{O}(\infty)$, $\mathrm{U}(p,\infty)/\mathrm{U}(p)\times\mathrm{U}(\infty)$, and $\mathrm{Sp}(p,\infty)/\mathrm{Sp}(p)\times\mathrm{Sp}(\infty)$, and that their isometry groups are, up to a subgroup of index two in the complex case, exactly the projective versions of $\mathrm{O}(p,\infty)$, $\mathrm{U}(p,\infty)$, and $\mathrm{Sp}(p,\infty)$ ([undefk], Theorem 3.3). Therefore, studying representations of $G$ that preserve a nondegenerate sesquilinear form of finite index is the same as studying actions of $G$ by isometries on infinite-dimensional symmetric spaces of nonpositive curvature and finite rank.
@@
In another paper [undefag], we consider the case where $G$ is a higher-rank algebraic group over a non-archimedean local field $\mathbb{F}$, and we show that, at least in the case $G=\mathrm{SL}_{n}(\mathbb{F})$ ($n\geq 3$), there are no continuous fixed point-free actions of $G$ on any infinite-dimensional symmetric space of nonpositive curvature and finite rank, which implies that there are no continuous irreducible representations of $G$ into any Pontryagin space.
Turning back to our paper, we will actually show Theorem 1.1 in more generality: not only for $G=\mathrm{Aut}(T)$, but for all closed noncompact subgroups of $\mathrm{Aut}(T)$ that act transitively on the boundary at infinity $\partial T$ and satisfy Tits’ independence property (P) ([undefaf], 4.2). We may also drop the assumption that the tree is regular and start from any tree where every vertex has degree $\geq 3$: our assumptions on $G$ imply that the tree is either regular or biregular (the latter meaning that there exist $r,s\geq 3$ such that all vertices at even distance from a given vertex have degree $r$, and all those at odd distance have degree $s$).
The groups we consider include universal groups associated with $2$-transitive permutation groups, which were introduced by Burger and Mozes in ([undeff], 3.2) and then studied by Amann in his PhD thesis ([undef], Chapter 3). More recently, Reid and Smith [undefac] provided a characterization of tree automorphism groups with Tits’ independence property (P) in terms of combinatorial structures called local action diagrams.
Unitary representations of the groups we consider have been classified by Amann ([undef], Chapter 2). The representations fall into three families: super-cuspidal, special, and spherical, according to the conjugacy class of a maximal compact subgroup that fixes a nontrivial vector.
Our result, formulated in this generality, is the following:
Theorem 1.2.
Let $T$ be a locally finite tree where every vertex has degree $\geq 3$. Let $G$ be a closed noncompact subgroup of $\mathrm{Aut}(T)$, with Tits’ Property (P) and transitive action on $\partial T$. Let $\mathbb{K}\in\{\mathbb{R},\mathbb{C},\mathbb{H}\}$, and let $(\pi,H)$ be a continuous irreducible representation of $G$ to a $\mathbb{K}$-Hilbert space. If $\pi$ preserves a continuous strongly nondegenerate sesquilinear form of finite index $p\geq 0$, then $p=0$ or $p=1$. Moreover, if $p=1$, then the representation $(\pi,H)$ is isomorphic to the tensor product with $\mathbb{K}$ of a real representation.
Here, $\mathbb{H}$ denotes the quaternions. Theorem 1.2 obviously implies Theorem 1.1, and the final statement about the case $p=1$ confirms that, also in our generality, the representations classified in [undefe] exhaust all possible non-unitary finite-index representations.
@@
The paper is organized as follows. In Section 2 we give a few preliminaries on trees and on the independence property, and we explain why we can restrict to a representation that is irreducible in the algebraic sense (the relevant technical results are shown in the Appendix). In Section 3 we show that also in our case we can classify the representations as spherical, special, and super-cuspidal, and that for our purposes it suffices to consider spherical representations. In Section 4 we consider the case where $G$ acts transitively on the set of vertices of the tree and give an explicit description of spherical representations, getting a classification that depends on only one parameter. In Section 5 we study nondegenerate sesquilinear forms of finite index that can be preserved by these representations, completing the proof of our theorem in the case where the action of $G$ is transitive on the vertices. Finally, in Section 6 we generalize the results of Sections 4 and 5 to the case where the action of $G$ is not transitive on the vertices. The Appendix, while functional to our scope of restricting to algebraically irreducible representations, contains results about admissibility of representations and irreducible actions of operators algebras that can also carry some independent interest.
@@ 2 Preliminaries
A tree $T$ is an undirected graph in which every two vertices are connected by exactly one path. We call the set of vertices $V$ and the set of (unoriented) edges $E$. A tree is regular if each vertex has the same degree $d$, i. e. it is an endpoint of exactly $d$ edges. This number $d$ is called the degree of the tree. A tree is biregular of degrees $(r,s)$ if all vertices at even distance from a base vertex have degree $r$, and all vertices at odd distance have degree $s$.
@@
We consider a locally finite tree $T$ where every vertex has degree $\geq 3$, and its automorphism group $\mathrm{Aut}(T)$. This group, equipped with the topology of pointwise convergence, is totally disconnected and locally compact. It has a natural action on the tree and on its boundary at infinity.
@@
We can now introduce the independence property for a subgroup $G<\mathrm{Aut}(T)$. We use the definition given by Amann ([undef], Definition 9):
@@
As observed by Amann (as a consequence of Lemma 10 in [undef]), if $G$ is a closed subgroup of $\mathrm{Aut}(T)$, this property is equivalent to Tits’ Property (P), which was defined in ([undefaf], 4.2). Therefore, from now on, we shall use Definition 2.1 as an equivalent definition for Tits’ Property (P).
In order to prove Theorem 1.2, we are interested in studying representations of closed noncompact subgroups of $\mathrm{Aut}(T)$ that act transitively on the boundary $\partial T$ and have the independence property. From now on, let $G$ be such a subgroup. We recall a few properties of $G$ that follow from the assumptions; more details can be found in [undef] (Proposition 4).
@@
    Either the action of $G$ on $V$ is transitive or it has two orbits. In the first case, the tree is regular; in the second case, the tree is regular or biregular, one orbit is the set of vertices at even distance from a fixed vertex $o$, and the other orbit is the set of vertices at odd distance from $o$.
When studying irreducible representations of $G$ into a Hilbert space $H$, irreducibility is intended in the topological sense: there are no closed $G$-invariant subspaces of $H$ except $\{0\}$ and $H$. We shall now reduce to a situation where we have irreducibility in the algebraic sense, i. e no $G$-invariant subspaces at all, except $\{0\}$ and the full space.
To this end, we restrict our attention to a dense subspace of $H$, as in the following proposition.
@@
Let $(\pi,H)$ be a continuous irreducible representation of $G$. Let $H^{(\infty)}\subseteq H$ be the subspace of smooth vectors in $H$, i. e vectors that are fixed by some open compact subgroup of $G$. Then $H^{(\infty)}$ is dense in $H$.
@@
Since the representation is irreducible and the subspace $H^{(\infty)}$ is $G$-invariant (if $g\in G$ and $K<G$ satisfies $Kv=v$, then $gv$ is fixed by the subgroup $gKg^{-1}$), it is enough to show that there exists one nonzero smooth vector.
Let $v\in H$, $v\neq 0$. By Hahn-Banach, there exists a continuous linear operator $L$ on $H$ such that $Lv\neq 0$. By the continuity of the representation, there is an open compact $K<G$ such that $|Lkv-Lv|<\frac{1}{2}|Lv|$ for all $k\in K$. We consider the vector $w:=\int_{K}kv\ dk$, where we integrate with respect to the Haar measure on $K$ (normalized so that the total measure of $K$ is $1$). This vector is fixed by $K$ and satisfies $|Lw-Lv|<\frac{1}{2}|Lv|$, which implies $|Lw|>\frac{1}{2}|Lv|$ and thus $w\neq 0$.
We will prove in the Appendix that, under the hypotheses of Theorem 1.2 (namely the existence of a preserved sesquilinear form of finite index), the representation $(\pi,H)$ is admissible, meaning that for every open compact $K<G$ the space $H^{K}$ of $K$-fixed vectors is finite-dimensional. This is not true in complete generality, as shown in an example by Nicolas Monod [undeft]. We will also show in the Appendix that admissibility implies that the restriction of the representation to $H^{(\infty)}$ is irreducible in the algebraic sense; therefore from now on we shall assume that we are dealing with an irreducible representation in the algebraic sense.
@@ 3 Restricting to spherical representations
In this section, we begin the classification of the representations. We will work on a group $G$ that satisfies the hypotheses of Theorem 1.2, and we will denote by $(\pi,H)$ a continuous algebraically irreducible representation of $G$ to a $\mathbb{K}$-vector space $H$ that is the subspace of smooth vectors of a $\mathbb{K}$-Hilbert space $\widetilde{H}$, with the topology inherited from $\widetilde{H}$. Here $\mathbb{K}\in\{\mathbb{R},\mathbb{C},\mathbb{H}\}$. In the case $\mathbb{K}=\mathbb{H}$, we will always intend $\mathbb{K}$-vector spaces as right vector spaces. We will also always assume all sesquilinear forms to be continuous.
@@
Given a continuous representation $(\pi,H)$ of $G$, we know from Proposition 2.3 that there is some open compact $K<G$ such that $H^{K}\neq\{0\}$, where the notation indicates the subspace of $K$-fixed vectors in $H$. We can consider the set of complete subtrees $S$ such that $H^{G(S)}\neq\{0\}$. What we have said, together with Proposition 2.2(a), implies that this set is not empty. Thus, we can consider the set $M$ of minimal complete subtrees $S$ in the set and distinguish three cases:
@@
[…]
@@
is a well-defined map from $H$ to $C(S)$: the function $g\mapsto\left\langle\xi,\pi(g^{-1})v\right\rangle$ is right-invariant under $G(S)$ because $\xi$ is invariant under $G(S)$, and left-invariant under some compact open subgroup of $G$ because every $v\in H$ is smooth. It is immediate to see that the map is $G$-equivariant and nonzero. Therefore, since the representation $(\pi,H)$ is irreducible, the map is injective and $(\pi,H)$ is a subrepresentation of $(\Pi_{S},C(S))$.
Amann proves ([undef], Lemma 20) that if the representation $(\pi,H)$ is super-cuspidal, the space $H\subset C(S)$ that arises by applying Proposition 3.2 to $S\in M$ consists of functions that are finitely supported when projected to $G/S$. This is shown for complex representations, but the proof is the same for real and quaternionic ones. Amann then proceeds to classify unitary super-cuspidal representations, but the proof that the functions are finitely supported on $G/S$ works without assuming that the representation is unitary. This follows as a consequence, as the representation will preserve the scalar product on $H$ defined as the restriction to $H$ of the standard scalar product on the space of finitely supported functions on $G/S$.
Similarly, Amann proves ([undef], Lemma 34) that if the representation $(\pi,H)$ is special, all functions in $H\subset C(S)$ (with $S$ consisting of two vertices linked by an edge) belong to $\ell^{2}(G/S)$. Again, the proof does not need that the representation is unitary, and this fact follows as a consequence, as the restriction to $H$ of the standard scalar product on $\ell^{2}(G/S)$ will be preserved.
In the rest of the paper, we will focus our attention on the remaining case, the case of spherical representations. Since in the other two cases all irreducible representations are unitary and since they satisfy $\dim(H^{G(S)})<\infty$ ([undef], Lemma 20, Lemma 34), they cannot preserve a nondegenerate sesquilinear form of index $\geq 1$. This is explained in the following proposition.
@@
Let $(\pi,H)$ be an irreducible $\mathbb{K}$-representation of $G$. Assume there exists an open compact subgroup $K$ such that $0<\dim(H^{K})<\infty$. If $\pi$ preserves a nondegenerate sesquilinear form $\langle\cdot,\cdot\rangle$ of index $p\geq 0$, with $2p\neq\dim(H)$, then any other sesquilinear form $\langle\langle\cdot,\cdot\rangle\rangle$ that it preserves is a real scalar multiple of $\langle\cdot,\cdot\rangle$.
Proof.
Any sesquilinear form $\langle\langle\cdot,\cdot\rangle\rangle$ on $H$ can be written as $\langle\langle\cdot,\cdot\rangle\rangle=\langle A\cdot,\cdot\rangle$, where $A:H\rightarrow H$ is a self-adjoint linear operator. If the form $\langle\langle\cdot,\cdot\rangle\rangle$ is preserved by the representation $\pi$, we have
for all $g\in G,\ v,w\in H$. This means that $A$ commutes with the representation, i. e. $A\in\mathrm{End}(\pi)$. By the Schur Lemma ([undefb], Lemma 3.21), $\mathrm{End}(\pi)$ is a division algebra over $\mathbb{R}$. If we fix a nonzero $v\in H^{K}$ the map $\mathrm{End}(\pi)\rightarrow H^{K}$ defined by $\phi\mapsto\phi(v)$ is injective, since $\pi$ is irreducible, thus $\dim(\mathrm{End}(\pi))\leq\dim(H^{K})$. Since $\dim(H^{K})$ is finite, so is $\dim(\mathrm{End}(\pi))$, and it is well known that then $\mathrm{End}(\pi)$ must be isomorphic to $\mathbb{R}$, $\mathbb{C}$ or $\mathbb{H}$.
If $\mathrm{End}(\pi)$ is isomorphic to $\mathbb{R}$, then we are done. In the other two cases, we have $A=\alpha Id+\beta A_{1}$ with $\alpha,\beta\in\mathbb{R}$ and $A_{1}^{2}=-Id$. We wish to prove that $\beta=0$. Since $A$ and $\alpha Id$ are self-adjoint and belong to $\mathrm{End}(\pi)$, the same must be true for $\beta A_{1}$ and therefore, if $\beta\neq 0$, also for $A_{1}$. Then for every $v\in H$ we have
Therefore, $A_{1}$ sends positive vectors (with respect to the form $\langle\cdot,\cdot\rangle$) to negative vectors and vice versa. As $A_{1}$ is invertible (because $A_{1}^{2}=-Id$), this means that it establishes a bijection between positive and negative vectors of $H$, which can only happen if the index $p$ of the form $\langle\cdot,\cdot\rangle$ satisfies $2p=\dim(H)$.
4 An explicit form for spherical representations
The space $C(S)$ is the space of $K$-right-invariant functions on $G$ that are left-invariant under some open compact subgroup. We can identify this space with the space of functions on $G/K$ (the set of right cosets of $G$ under $K$) that are left-invariant under some compact open subgroup of $G$. The set $G/K$ can be identified with the orbit of $o$ in $T$ under the action of $G$, which is either the full vertex set $V$ or the set of vertices at even distance from $o$ (Proposition 2.2(c)).
In this section, we assume that we have the full vertex set $V$, i. e. that the action of $G$ is transitive on the vertices. In this case, the tree must be regular. In Section 6 we will extend the results to the other case.
Proposition 3.2 tells us that we can see the representation $(\pi,H)$ as a subrepresentation of $(\Pi,C)$, where $C$ is the space of $\mathbb{K}$-valued functions on $V$ that are left-invariant under some open compact subgroup of $G$ and $\Pi$ is the left action of $G$ on $C$. Therefore, from now on, we assume that $H\subset C$ and $\pi$ is the restriction of $\Pi$ to $H$.